OpenStudy (abbles):

Le can clean the house in 9 hr, while Tran needs 6 hr. How long will it take them to clean the house if they work together?

OpenStudy (abbles):

@Rishi123 @kropot72 Do you guys know?

OpenStudy (kropot72):

In one hour Le working alone can clean 1/9 of the house, and in one hour Tran working alone can clean 1/6 of the house. Therefore working together for one hour, Le and Tran can clean: (1/9) + (1/6) of the house. @Abbles Can you follow so far?

OpenStudy (kropot72):

@Abbles Are you there?

OpenStudy (abbles):

Yes... so where would I go from there?

OpenStudy (abbles):

(1/9) + (1/6) = x ? Or maybe (1/9) + (1/6) = (1/x) ?

OpenStudy (kropot72):

\[\large \frac{1}{9}+\frac{1}{6}=\frac{5}{18}\] Therefore in one hour, working together, Le and Tran can clean 5/18 of the house. So in 1/5 of an hour they can clean 1/18 of the house, meaning that they must work 18/5 hours together to clean the whole house.

OpenStudy (abbles):

Oh my gosh you are brilliant. Thank you so much! That cleared it up big time.

OpenStudy (abbles):

So together it takes them 3.6 hours. Thank you very much

OpenStudy (kropot72):

You're welcome :)